# Classification of curl forces for all space dimensions

### v2

@@ Abstract
We present a decomposition of classical potentials into a conservative (gradient) component and a non-conservative component. The latter generalizes the curl component of the force in the three-dimensional case. The force is transformed into a differential $1$-form, known as the work form. This work form is decomposed into an exact (gradient) component and an antiexact component, which in turn generalizes the curl part of the force. The antiexact component is subsequently decomposed using the Frobenius theorem. This local decomposition is a useful tool for identifying the specific components of classical potentials.
@@ 1 Introduction
A well-known result in classical mechanics [24, 1, 12] is that a force with a vanishing curl is the gradient of a scalar function known as a potential. On a Riemannian manifold [2, 8, 20, 24, 15, 22, 1, 12, 23] $M$ with a metric tensor $g:TM\times TM\rightarrow\mathbb{R}$, a force can be represented by a differential $1$-form (the work form). In this framework, the result is equivalent to the statement that the work form is exact, meaning it is the exterior derivative of a 0-form (a scalar function), which is the potential. This property holds locally by the Poincaré lemma [10, 11, 21].
A force $F$ satisfying the condition $\nabla\times F\neq 0$ is known as a curl force, also referred to as a pseudo-gyroscopic or circulatory force. Such a force can still be decomposed into gradient-like components, although this decomposition is not unique. These forces are of physical significance [3, 5, 6] and can be quantized [7]. The classical motions under these forces can be classified [4]. They also imply also some topological properties of magnetic fields [16]. Therefore, understanding their structure is of paramount importance.
In [26] a decomposition of a curl force expressible by functions and their gradients was given, namely:
@@
  • Proposition 2.4 [26]: In $\mathbb{R}^{3}$ a force $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ can be expressed as $F(x)=-V(x)\nabla U(x)-\nabla W(x)$, where generalized potentials $U,V,W:\mathbb{R}^{2}\rightarrow\mathbb{R}$, and $\nabla V\times\nabla U\neq 0$. Moreover, $\nabla V(x)=0\Leftrightarrow\nabla\times F=0$ and we can select $W=0$. We have $\nabla W=0\Leftrightarrow F\cdot\nabla\times F=0$.
These results stem from transforming [^1] The mappings $g:TM\rightarrow T^{*}M$ and $g^{-1}:T^{*}M\rightarrow TM$ are called musical isomorphisms [24, 1, 12]. a force $F\in TM$ into the work form $\omega=g(F,\_):TM\rightarrow\mathbb{R}$, i.e., $\omega\in T^{*}M$. Applying the Darboux theorem [10, 11] then determines the local form of $\omega$, which is subsequently transformed back into a vector field on the $TM$ bundle. Due to the use of the Darboux theorem, the resulting decomposition is local in nature and not unique. This approach is therefore distinct from the global Helmholtz-Hodge decomposition [20, 25], which is valid for compact manifolds. Moreover, these findings suggest that it is more natural to work with forces by expressing them as differential forms, particularly as the standard notion of a curl does not exist in higher-dimensional spaces.
In this paper, we propose a new, algorithmic approach to decompose any force into a potential component and a generalized curl component. The curl component is then non-uniquely decomposed into terms involving generalized potentials and their gradients using the Frobenius theorem. Unlike the method outlined in [26], our proposed decomposition does not require solving PDEs. This method provides a practical way of working with classical potentials in any number of dimensions. To address the difficulty of defining a curl outside of three-dimensional space, we replace it with the more general notion of an antiexact differential form. The proposed method also yields a more fine-grained decomposition than the one proposed in [26].
The paper is organized as follows. In the next section, we summarize the key results from the theory of exterior differential forms that are needed to construct the decomposition algorithm. We then present our main result and provide several examples that illustrate the algorithm’s application.
@@ 2 Preliminaries
In this section we summarize results from local behaviour of exterior differential forms following [10, 11, 21, 17, 18, 14]. They will be used for the construction of the force decomposition in the next section.
The main object of consideration is an $n$-dimensional smooth manifold $M$ equipped with a metric tensor $g:M\rightarrow\mathbb{R}$, since an inner product is defined in most physical problems. We will also use the exterior derivative $d:\Lambda^{k}(M)\rightarrow\Lambda^{k+1}(M)$.
On this manifold, we can associate a vector field $F\in TU$ with a $1$-form $\omega$ musical isomorphism $\omega=g(F,\_)$. Similarly, the gradient of a function $f:U\rightarrow\mathbb{R}$ defined via its differential $df$ such that $df=g(\nabla f,\_)$.
The curl operation is defined only in three-dimensional space. This is because the cross product, and consequently the curl, is specific to three-dimensional Euclidean space, where an oriented plane (a bivector) can be uniquely represented by a vector [9, 12]. This representation is difficult in curvilinear coordinates and not possible in higher-dimensional spaces. A good summary of this situation is the statement from [12]: ”To my knowledge, however, there is no reason for ever writing out the components of the vector curl A in curvilinear coordinates; if the expression curl A appears, it is a sure sign that the vector in question was not the contravariant A but rather the covariant vector (…)”. Therefore, to operate in spaces of any dimension, one must abandon the notion of the curl and switch to the more general notion of the exterior derivative acting on differential forms. We will see that the natural generalization of the curl component of a vector is an antiexact differential form, which we define below. We also must note that the Helmholtz decomposition can be constructed in vector calculus [13], where the construction of multivectors is needed.
@@ 2.1 Geometric decomposition
We restrict our analysis to a star-shaped subset of $M$ that lies within a single local chart, which can therefore be identified with a star-shaped subset $U\in\mathbb{R}^{n}$. The metric tensor $g:M\rightarrow\mathbb{R}$ can then be restricted to $U$ in a chart.
The star-shaped domain $U$ contains a central point $x_{0}\in U$ and admits a linear homotopy $F:[0,1]\times U\rightarrow U$, $F(t,x)=x_{0}+t(x-x_{0})$ which connects any point $x\in U$ with $x_{0}$. We define the linear homotopy operator as
@@
where $\mathcal{K}=(x-x_{0})^{i}\partial_{i}$, $\omega\in\Lambda^{k}(U)$ is a $k$-differential form on $U$, and $i:\Lambda^{k}(U)\rightarrow\Lambda^{k-1}(U)$ is the insertion operator $i_{\mathcal{K}}\omega=\omega(\mathcal{K},\ldots)$. One can use a different homotopy in the definition, however the linear one makes the operator nilpotent $H^{2}=0$. It has also properties $dHd=d$ and $HdH=H$. We have also homotopy invariance formula
@@
where $s_{x_{0}}^{*}$ is the pullback along the constant map $s_{x_{0}}:x_{0}\hookrightarrow U$. The map $s_{x_{0}}^{*}$ can be nonzero only on $\Lambda^{0}(U)$. Since we will be considering $1$-forms, the pullback $s_{x_{0}}^{*}$ vanishes, and from (2) it follows that $Hd+dH=I$ on $\Lambda^{k}(U)$ for $k>0$. Moreover, since $(dH)^{2}=dH$ and $(Hd)^{2}=Hd$, the operators $Hd$ and $dH$ are projectors, and they define a direct sum decomposition of $\Lambda(U)$ into the exact [^2] Since the domain is star-shaped, an exact from is also closed (i.e., an elements of $ker(d)$) by the Poincare lemma. vector space $\mathcal{E}^{k}(U)=ker(d)=im(dH)$, and the antiexact module over $\Lambda^{0}(U)=C^{\infty}(U)$ denoted by $\mathcal{A}^{k}(U)=ker(H)=im(Hd)=\{\omega\in\Lambda^{k}(U)|H\omega=0,\omega|_{x_{0}}=0\}$. For $f\in\Lambda^{0}(U)$ we also have decomposition [^3] If $f$ is a classical potential then the decomposition agrees with the global gauge freedom realized by the additive group $(\mathbb{R}=ker(d),+)$. $f(x)=\underbrace{(f(x)-f(x_{0})}_{\mathcal{A}}\oplus\underbrace{f(x_{0})}_{\mathcal{E}}$ as $f(x_{0})\in ker(d)$, $f(x)-f(x_{0})\in ker(H)$ and $(f(x)-f(x_{0}))|_{x_{0}}=0$. In summary, these statements yield the decomposition
@@
The Poincare lemma can be proved using the geometric decomposition as $d\omega=0\Leftrightarrow\omega=dH\omega$, that is, if $\omega$ is closed, then it is also exact on a star-shaped subset.
@@ 2.2 Frobenius theorem
and treat them as a system of exterior differential equations. The form $\Sigma$ is called the torsion [^4] This is not the torsion from the Cartan moving frame method on Riemannian manifolds [22]. It describes only an obstruction to complete integrability of $\Omega$ with analogy to Riemannian curvature $2$-form for integrability of geodesics. It is standard equation of fundamental theories of physics - it is visible when it is rewritten as $d\Omega-\Gamma\wedge\Omega=\Sigma$ and identify the operator $d-\Gamma\wedge$\_ as a simplified version of exterior covariant derivative, see [19]. of the equation. The solution of this system is given by the following
@@
These results do not change if $\Omega$ is $k$-form for $k>1$; see [10] Section 5. One can also generalize these results to vector-valued differential forms; see [10, 19].
3 Force decomposition
We now combine geometric decomposition and the Frobenius theorem to decompose the work form into simpler components.
We begin by considering the work $1$-form $\omega$ corresponding to a force field $F:M\rightarrow TM$, defined as $\omega=g(F,\_)$. Since our method uses the homotopy operator, we restrict our analysis to a single chart, working within a star-shaped subset $U\subset\mathbb{R}^{n}$.
@@ 3 Force decomposition
    $h=H(e^{-g}\Omega),~\eta=H(e^{-g}\Sigma^{\prime})=Hd(e^{-g}\Omega).$ (19)
@@
    $\Omega=e^{g}(dh+H(\theta\wedge dh)),$ (20)
@@
    $\Omega=e^{g}dh,~h=H(e^{-g}\Omega).$ (21)
The functions $f,g,h$ can be called generalized potentials. The function $f$ corresponds to the standard potential when $\Omega=0$ when $g=0=h$.
@@
Proof.
$\omega=dH\omega+Hd\omega.$ (22)
We identify in the exact (’gradient’) part the potential $f=H\omega\in\Lambda^{0}(U)$. The second term, $\Omega=Hd\omega\in\Lambda^{1}(U)$, is the antiexact (’curl’) part as.
@@
  • General case: $X=e^{g}\nabla h+Y$, where $Y=e^{g}g^{-1}(\eta,\_)$.
  • Recursive case: $X=\nabla h+e^{g}Y$, where $Y=g^{-1}(H(\theta\wedge dh),\_)$.
  • Gradient recursive case: $X=e^{g}\nabla h$.
@@
Input: A force $F$, a star-shaped subset $U\in\mathbb{R}^{n}$, and a metric tensor $g$.
@@
  2. 2.
    Compute the exact part $dH\omega$ of $\omega$ and infer the potential $f=H\omega$.
@@
    Compute the antiexact part $\Omega=Hd\omega$.
@@
      Compute $d\Omega=dHd\Omega=d\omega$.
    2. (b)
      Select forms $\Gamma$ and $\Sigma$ that satisfy the equation $d\Omega=\Gamma\wedge\Omega+\Sigma$ (see Remark 4 on arbitrariness of this choice).
    3. (c)
      Compute $g=H\Gamma$, $h=H(e^{-g}\Omega)$, and $\eta=Hd(e^{-g}\Omega)$.
    4. (d)
      Construct the final decomposition: $\omega=df+e^{g}(dh+\eta)$.
  5. 5.
@@
The seemingly arbitrary selection of $\Gamma$ and $\Sigma$ for the case $\Omega\neq 0$ can be simplified in practice. As shown in Example 2 below, one can first select the factor $e^{-g}$, which fixes $\Gamma=dg$. Then, one can compute $h=H(e^{-g}\Omega)$ and $\eta=Hd(e^{-g}\Omega)$. In this approach, it is not necessary to compute $\Sigma$ explicitly.
We now provide an example of the application of this decomposition.

### v3

# Decomposition and characterization of curl forces for all space dimensions

@@ Abstract
This paper introduces a PDE-free algorithmic framework for the local decomposition of classical forces in arbitrary dimensions. By representing a force field as a differential $1$-form (work form), we employ the homotopy operator on a star-shaped domain to achieve a geometric decomposition into exact (gradient) and antiexact components. The antiexact part serves as a formal generalization of the curl force - or circulatory force - outside of three-dimensional Euclidean space. To further characterize the non-conservative dynamics, we apply the Frobenius theorem to the antiexact component, resolving it into integrable terms associated with generalized potentials and a path-dependent ’core’ representing fundamental obstructions to integrability. Unlike the Darboux-based classification, this constructive approach bypasses the requirement for solving partial differential equations, offering a practical tool for analyzing non-autonomous influences and scaling effects in complex physical systems.
@@ 1 Introduction
A well-known result in classical mechanics [24, 1, 12] is that a force with a vanishing curl is the gradient of a scalar function known as a potential. On a Riemannian manifold [2, 8, 20, 24, 15, 22, 1, 12, 23] $M$ with a metric tensor $g:TM\times TM\rightarrow\mathbb{R}$, a force can be represented by a differential $1$-form (the work form). In this framework, the existence of potential is equivalent to the statement that the work form is exact, meaning it is the exterior derivative of a 0-form (a scalar function), which is the potential. This property holds locally by the Poincaré lemma [10, 11, 21].
A force $F$ satisfying the condition $\nabla\times F\neq 0$ is known as a curl force, also referred to as a pseudo-gyroscopic or circulatory force. Such a force can still be decomposed into gradient-like components, although this decomposition is not unique. These forces are of physical significance [3, 5, 6] and some particular cases of them can be quantized [7]. The classical motions under these forces can be classified [4]. They also imply some topological properties of magnetic fields [16]. Therefore, understanding their structure is of paramount importance.
A decomposition of a curl force expressible by functions and their gradients was given in [26], namely:
@@
  • Proposition 2.4 [26]: In $\mathbb{R}^{3}$ a force $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ can be expressed as $F(x)=-V(x)\nabla U(x)-\nabla W(x)$, where generalized potentials are $U,V,W:\mathbb{R}^{2}\rightarrow\mathbb{R}$, and $\nabla V\times\nabla U\neq 0$. Moreover, $\nabla V(x)=0\Leftrightarrow\nabla\times F=0$ and we can select $W=0$. We have $\nabla W=0\Leftrightarrow F\cdot\nabla\times F=0$.
These results stem from transforming [^1] The mappings $g:TM\rightarrow T^{*}M$ and $g^{-1}:T^{*}M\rightarrow TM$ are called musical isomorphisms [24, 1, 12]. a force $F\in TM$ into the work form $\omega=g(F,\_):TM\rightarrow\mathbb{R}$, i.e., $\omega\in T^{*}M$; applying the Darboux theorem [10, 11] then determines the local form of $\omega$, which is subsequently transformed back into a vector field on the $TM$ bundle. Due to the use of the Darboux theorem, the resulting decomposition is local in nature and not unique. This approach is therefore distinct from the global Helmholtz-Hodge decomposition [20, 25], which is valid for compact manifolds. Moreover, these findings suggest that it is more natural to work with forces by expressing them as differential forms, particularly as the standard notion of a curl does not exist in higher-dimensional spaces.
In this paper, we propose a new, algorithmic approach to decompose any force into a potential component and a generalized curl component. This will be done using geometric decomposition. The curl component is then non-uniquely decomposed into terms involving generalized potentials and their gradients using the Frobenius theorem. Unlike the method outlined in [26], our proposed decomposition does not require solving PDEs. This PDE-free alternative, unlike the Darboux theorem, offers a novel and practical way of working with classical potentials in any number of dimensions. To address the difficulty of defining a curl outside of three-dimensional space, we replace it with the more general notion of an antiexact differential form. The proposed method also yields a more fine-grained decomposition than the one proposed in [26].
The paper is organized as follows: In the next section, we summarize the key results from the theory of exterior differential forms that are needed to construct the decomposition algorithm. We then present our main result and provide several examples that illustrate the algorithm’s application.
@@ 2 Preliminaries
In this section, we summarize results from the local behavior of exterior differential forms following [10, 11, 21, 17, 18, 14]. They will be used to construct the force decomposition in the next section.
The main object of consideration is an $n$-dimensional smooth manifold $M$ equipped with a metric tensor $g:TM\times TM\rightarrow\mathbb{R}$, since an inner product is defined in most physical problems. We will also use the exterior derivative $d:\Lambda^{k}(M)\rightarrow\Lambda^{k+1}(M)$.
On this manifold, we can associate a vector field $F\in TU$ with a $1$-form $\omega$, using musical isomorphism $\omega=g(F,\_)$. Similarly, the gradient of a function $f:U\rightarrow\mathbb{R}$ is defined via its differential $df$ such that $df=g(\nabla f,\_)$.
The curl operation is defined only in three-dimensional space. This is because the cross product, and consequently the curl, is specific to three-dimensional Euclidean space, where an oriented plane (a bivector) can be uniquely represented by a vector [9, 12]. This representation is difficult in curvilinear coordinates and not possible in higher-dimensional spaces. Therefore, to operate in spaces of arbitrary dimension, one must abandon the notion of the curl and switch to the more general notion of the exterior derivative acting on differential forms. We will see that the natural generalization of the curl component of a vector is an antiexact differential form, as defined below.
We also note that the Helmholtz decomposition can be constructed in vector calculus [13], where multivectors are needed.
@@ 2.1 Geometric decomposition
We restrict our analysis to a star-shaped subset of $M$ that lies within a single local chart, which can therefore be identified with a star-shaped subset $U\in\mathbb{R}^{n}$. The metric tensor $g:TM\times TM\rightarrow\mathbb{R}$ can then be restricted to $U$ in a chart.
The star-shaped domain $U$ contains a central point $x_{0}\in U$ and admits a linear homotopy $F:[0,1]\times U\rightarrow U$, $F(t,x)=x_{0}+t(x-x_{0})$, which connects any point $x\in U$ with $x_{0}$. We define the linear homotopy operator as
@@
where $\mathcal{K}=(x-x_{0})^{i}\partial_{i}$, $\omega\in\Lambda^{k}(U)$ is a $k$-differential form on $U$, and $i:\Lambda^{k}(U)\times TU\rightarrow\Lambda^{k-1}(U)$ is the insertion operator $i_{\mathcal{K}}\omega=\omega(\mathcal{K},\ldots)$. One can use a different, e.g., nonlinear, homotopy $F$ in the definition of $H$, however the linear one makes the operator nilpotent $H^{2}=0$. It also has the properties $dHd=d$ and $HdH=H$. We also have a homotopy invariance formula
@@
where $s_{x_{0}}^{*}$ is the pullback along the constant map $s_{x_{0}}:x_{0}\hookrightarrow U$. The map $s_{x_{0}}^{*}$ can be nonzero only on $\Lambda^{0}(U)$. Since we will be considering $1$-forms, the pullback $s_{x_{0}}^{*}$ vanishes, and from (2) it follows that $Hd+dH=I$ on $\Lambda^{k}(U)$ for $k>0$. Moreover, since $(dH)^{2}=dH$ and $(Hd)^{2}=Hd$, the operators $Hd$ and $dH$ are projectors. They define a direct sum decomposition of $\Lambda(U)$ into the exact [^2] Since the domain is star-shaped, an exact form is also closed (i.e., an element of $ker(d)$) by the Poincaré lemma. vector space $\mathcal{E}^{k}(U)=ker(d)=im(dH)$, and the antiexact module over $\Lambda^{0}(U)=C^{\infty}(U)$ denoted by $\mathcal{A}^{k}(U)=ker(H)=im(Hd)=\{\omega\in\Lambda^{k}(U)|H\omega=0,\omega|_{x_{0}}=0\}$. For $f\in\Lambda^{0}(U)$, we also have a decomposition [^3] If $f$ is a classical potential, then the decomposition agrees with the global gauge freedom realized by the additive group $(\mathbb{R}=ker(d),+)$. $f(x)=\underbrace{(f(x)-f(x_{0})}_{\mathcal{A}}\oplus\underbrace{f(x_{0})}_{\mathcal{E}}$ as $f(x_{0})\in ker(d)$, $f(x)-f(x_{0})\in ker(H)$ and $(f(x)-f(x_{0}))|_{x_{0}}=0$. In summary, these statements yield the direct sum decomposition
@@
The Poincaré lemma can be proved using the geometric decomposition as $d\omega=0\Leftrightarrow\omega=dH\omega$, that is, if $\omega$ is closed, then it is also exact on a star-shaped subset.
@@ 2.2 Frobenius theorem
and treat them as a system of exterior differential equations. The form $\Sigma$ is called the torsion [^4] This is not the torsion from the Cartan moving frame method on Riemannian manifolds [22]. It describes only an obstruction to the complete integrability of $\Omega$ with analogy to the Riemannian curvature $2$-form for the integrability of geodesics. It is standard equation of fundamental theories of physics - it is visible when it is rewritten as $d\Omega-\Gamma\wedge\Omega=\Sigma$ and identify the operator $d-\Gamma\wedge$\_ as a simplified version of exterior covariant derivative, see [19]. of the equation. The solution of this system is given by the following
@@
These results do not change if $\Omega$ is a $k$-form for $k>1$; see [10] Section 5. One can also generalize these results to vector-valued differential forms; see [10, 19].
3 Results
We now combine geometric decomposition and the Frobenius theorem to decompose the work form into simpler components. Both of these methods are already known; however, the novelty of the paper lies in their subsequent application to obtain a PDE-free decomposition of arbitrary forces, in contrast to the Darboux-based approach.
We begin by considering the work $1$-form $\omega$ corresponding to a force field $F:M\rightarrow TM$, defined as $\omega=g(F,\_)$. Since our method uses the homotopy operator, we restrict our analysis to a single chart and work within a star-shaped subset $U\subset\mathbb{R}^{n}$.
@@ 3 Results
    $\phi=H(e^{-\gamma}\Omega),~\eta=H(e^{-\gamma}\Sigma^{\prime})=Hd(e^{-\gamma}\Omega).$ (19)
@@
    $\Omega=e^{\gamma}(d\phi+H(\theta\wedge d\phi)),$ (20)
@@
    $\Omega=e^{\gamma}d\phi,~h=H(e^{-\gamma}\Omega).$ (21)
The functions $f,\gamma,\phi$ can be called generalized potentials. The function $f$ corresponds to the standard potential when $\Omega=0$ when $\gamma=0=\phi$.
Proof.
$\omega=dH\omega+Hd\omega.$ (22)
We identify in the exact (’gradient’) part the potential $f=H\omega\in\Lambda^{0}(U)$. The second term, $\Omega=Hd\omega\in\Lambda^{1}(U)$, is the antiexact (’curl’) part.
@@
  • General case: $X=e^{\gamma}\nabla\phi+Y$, where $Y=e^{\gamma}g^{-1}(\eta,\_)$.
  • Recursive case: $X=\nabla\phi+e^{\gamma}Y$, where $Y=g^{-1}(H(\theta\wedge d\phi),\_)$.
  • Gradient recursive case: $X=e^{\gamma}\nabla\phi$.
@@
Input: A force $F$ and a metric tensor $g$.
@@
    Select the homotopy center $x_{0}$ and define a star-shaped subset $U\in\mathbb{R}^{n}$ for $x_{0}$.
  2. 2.
@@
    Extract the potential $f=H\omega$.
@@
    Identify the generalized curl component $\Omega=Hd\omega$.
  5. 5.
@@
      Apply the Frobenius decomposition to $\Omega$ based on the system’s integrability (General, Recursive, or Gradient Recursive).
  6. 6.
@@
The seemingly arbitrary selection of $\Gamma$ and $\Sigma$ for the case $\Omega\neq 0$ can be simplified in practice. As shown in Example 2 below, one can first select the factor $e^{-\gamma}$, which fixes $\Gamma=d\gamma$. Then, one can compute $\phi=H(e^{-\gamma}\Omega)$ and $\eta=Hd(e^{-\gamma}\Omega)$. In this approach, it is not necessary to compute $\Sigma$ explicitly.
We also summarize the proposed algorithm with the Darboux decomposition in Table 1.
Feature Darboux approach Algorithm 1
Computational Requirement Solving PDEs Integrations (Homotopy op. $H$)
Uniqueness Non-unique Non-unique (Gauge choice, homotopy center choice)
Dimensionality 2D/3D Arbitrary $n$-dimensions
Components Potential + Curl Potential + General/Recursive Curl(1-2 components)
Table 1: Comparison of the Darboux approach [26] and Algorithm 1.
3.1 Non-uniqueness considerations
Once the homotopy center $x_{0}$ is fixed, the antiexact part $\Omega$ is a geometric object that has a fixed transformation under the change of the center, see [10].
The choices of $\Gamma$ and $\Sigma$ is arbitrary as long as they fulfil $d\Omega=\Gamma\wedge\Omega+\Sigma$. The selection of $\Gamma$, and consequently $\gamma$, can be seen as a ’gauge’ choice - the integrating factor $e^{-\gamma}$ can be chosen to simplify decomposition. Different choices of this factor balance the complexity of the factors between the $e^{\gamma}d\phi$ and the path-dependent $e^{\gamma}\eta$.
The specific choice of $\Gamma$ can be predicted by a physical situation, e.g., by the preservation of symmetry. The other choice can be guided by ’minimizing’ the torsion $\Sigma$ to decrease the level of non-integrability of $\Omega$.
3.2 Examples
We now provide an example of its application.